# Weighted shift operators on spaces of analytic functions over an annulus - I

### v1

@@ Abstract.
We initiate a study of the forward weighted shift operator, denoted by $F_{w}$, defined on the Banach spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ of analytic functions on a suitable annulus $\mathtt{A}$ in the complex plane. Here, $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a.b}(\mathtt{A})$, respectively, are the spaces of analytic functions on some annulus $\mathtt{A}$, having normalized Schauder bases of the form $\{(a_{n}+b_{n}z)z^{n},~{}~{}n\in\mathbb{Z}\}$, which are equivalent to the standard basis in $\ell^{p}(\mathbb{Z})$ and $c_{0}(\mathbb{Z})$. We first obtain necessary and sufficient conditions for $F_{w}$ to be bounded, and then we primarily study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, supercyclic, and chaotic. It is proved that, under certain conditions on $\{a_{n}\}$ and $\{b_{n}\}$, the operator $F_{w}$ similar to a compact perturbation of a classical bilateral weighted shift on $\ell^{p}(\mathbb{Z})$. Unlike the case of classical weighted shifts, the adjoint operator $F_{w}^{*}$ on the dual of $\ell^{p}_{a,b}(\mathtt{A})$ can have non-trivial periodic vectors, without being even hypercyclic. The zero-one law of orbital limit points fails for $F_{w}^{*}$.
@@ 1. Introduction
Among various classes of operators that are commonly investigated in operator theory and linear dynamics, the class of weighted shifts has a predominant role. Shift operators are fairly simple to define, yet they yield diverse and appealing results. In operator theory, weighted shifts are well studied in connections with model theory (representation of various operators using shifts), invariant subspaces, cyclic vectors, subnormality, and so on, cf. Halmos [24], and Shields [38]. In the context of linear dynamics it is a usual practice that dynamical properties are first tested for weighted shifts defined on the sequence spaces $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$. Important dynamical notions such as hypercyclicity, chaos, and ergodicity are well distinguishable using weighted shifts. Motivated these, in this paper we introduce and study weighted shift operators on spaces $\ell^{p}_{a,b}(\mathbb{D})$ and $\ell^{p}_{a,b}(\mathtt{A})$ having a close resemblance to the weighted $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ spaces, highlighting the similarities and differences in the dynamics. These newly introduced spaces and weighted shifts are of independent interests as well.
@@
  • topologically transitive if, for two non-empty open sets $U_{1}$ and $U_{2}$ of $X$, there exists a natural number $k$ such that $T^{k}(U_{1})\cap U_{2}\neq\phi$.
@@
Rolewicz [35] showed that $\lambda B$ is hypercyclic on $\ell^{p}(\mathbb{N})$, $1\leq p<\infty$, where $|\lambda|>1$ and $B$ is the unweighted backward shift operator. Also, for particular classes of hypercyclic weighted shifts, see Kitai [25], Gethner and Shapiro [19], and Godefroy and Shapiro [20]. A complete characterization of hypercyclic weighted shifts on $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ was obtained by Salas [36]. In [20] the authors introduced and studied the notion of chaos for operators, and established that several familiar operators including weighted shifts are chaotic. A complete characterization of hypercyclic and chaotic weighted shifts in a general $F$-spaces was obtained by Grosse-Erdmann [22]. We also note that the weighted shifts which are (topologically) mixing were characterized by Costakis and Sambarino [13]. The dynamics of the unweighted backward shift on certain Bergman spaces was studied in [32]. For account on the theory on hypercyclicity and related topics, we refer to the monographs by Bayart and Matheron [3], and Grosse-Erdmann and Peris [23]. In practice, the existence of hypercyclic and chaotic properties of operators is established using well known criteria that are stated below, cf. [3] and [23]. We will make use of these standard criteria to obtain the hypercyclic and chaotic properties of weighted shifts defined on the weighted spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ which are introduced in Section $2$.
@@
(Gethner-Shapiro Criterion [19]) Let $T$ be a bounded operator on a separable Banach space $X$, and let $X_{0}$ be a dense subset of $X$. If $\{n_{k}\}\subseteq\mathbb{N}$ is a strictly increasing sequence and $S:X_{0}\mapsto X_{0}$ is a map such that, for each $x\in X_{0}$, $lim_{k\rightarrow\infty}T^{n_{k}}x=0=\lim_{k\rightarrow\infty}S^{n_{k}}x,$ and $TSx=x$, then $T$ is hypercyclic. in addition, if $n_{k}=k$ for all $k\geq 1$, then $T$ is mixing on $X$.
@@
(Chaoticity Criterion [9]) Let $X$ be a separable Banach space, $X_{0}$ be a dense set in $X$, and let $T$ be a bounded operator on $X$. If there exists a map $D:X_{0}\rightarrow X_{0}$ such that $\sum_{n\geq 0}T^{n}x$ and $\sum_{n\geq 0}S^{n}x$ are unconditionally convergent, and $TSx=x$, for each $x\in X_{0}$, then the operator $T$ is chaotic and mixing on $X$.
@@
(Supercyclicity Criterion [31]) For a bounded operator $T$ on a separable Banach space $X$, if there exist a dense subset $X_{0}$, a sequence $(n_{k})$ of natural numbers, and a map $S:X_{0}\rightarrow X_{0}$ such that, for all $x,y\in X_{0}$, $\lim_{k\rightarrow\infty}\|T^{n_{k}}x\|\|S^{k}y\|=0$ and $TSx=x$, then $T$ is supercyclic.
Motivated by the classical bilateral shifts, in this paper we initiate a study of weighted bilateral forward shift operators $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. In Section $2$, we introduce the aforementioned spaces as analytic function spaces and prove that the co-ordinate functionals on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ are continuous. (Although sequential analogues of these spaces can be defined, we prefer to stick to the analytic function space set up since we believe that our study will be useful in the operator theoretic aspects such as descriptions of commutants, cyclic vectors, and so on.) In Section $3$ we obtain necessary and sufficient conditions for the boundedness of $F_{w}$, and as main results, we study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, mixing, supercyclic, and chaotic. Dynamics of the adjoint of a unilateral forward shift is studied in Section $4$. In Section $5$, it is shown that the bilateral shifts on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ and also the unilateral shifts have distinguishing properties compared to classical weighted shifts . For example, among other results, it is shown that there are unilateral and bilateral forward shifts $F$ such that $F^{*}$ has a non-zero periodic vector, without being hypercyclic.
@@
if and only if $\lVert f\rVert_{\ell^{p}_{a,b}}:=\left(\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert^{p}\right)^{\frac{1}{p}}<\infty.$ Note that for $b_{n}=0$ for all $n\in\mathbb{Z},$ we get a space that is equivalent to the standard weighted $\ell^{p}(\mathbb{Z})$. In a similar way, we define the space $c_{0,a,b}(\mathtt{A})$ which consists of analytic functions, for which $\{(a_{n}+b_{n}z)z^{n}:n\in\mathbb{Z}\}$ forms a normalized Schauder basis, equivalent to the standard basis in $c_{0}(\mathbb{Z})$. Moreover, for $f=\sum_{n}\lambda_{n}f_{n}$ in $c_{0,a,b}(\mathtt{A})$, we have $\lambda_{n}\rightarrow 0$, as $|n|\rightarrow\infty$. Define its norm as
@@
A few comments are in order. The natural domains $\mathtt{A}_{j}$, $j=1,2,3$, will always be assumed to be non-degenerate. Secondly, in the case of the usual weighted spaces, the natural domains of the map $z\mapsto ev_{z}$ are open or closed annuli. However, this is not so for $\ell^{p}_{a,b}(\mathtt{A})$ which is an important property that distinguishes these spaces from the classical weighted $\ell^{p}$ spaces, which will be discussed and utilized in Section $5$. Also, the series $\sum_{n=-\infty}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}$ can still converge at a point $z$ that lies outside $A_{j}$, and hence the domain of $z\mapsto ev_{z}$ can be extended to “extra points” outside $\mathtt{A_{j}}$, cf. Section $5$. We note that such a phenomenon in the unilateral case when $p=2$, was observed by Adams and McGuire [1] who introduced the space the analytic reproducing kernel Hilbert space $\ell^{2}_{a,b}$ having an orthonormal basis of the form $(a_{n}+b_{n}z)z^{n}$, $n\geq 0$.
Since the evaluation functional $ev_{z}$ is continuous for all $z$, we now get the continuity of the co-ordinate functionals which are defined as follows. Since each $f(z)$ in $\ell^{p}_{a,b}(\mathtt{A})$ is analytic on the annulus $\mathtt{A}_{1}$, we have the Laurent expansion $f(z)=\sum_{n=-\infty}^{\infty}\widehat{f}(n)z^{n}$, $z\in\mathtt{A}_{1}$, where $\widehat{f}(n)$ is the $n$-th Laurent coefficient of $f(z)$. The $n$-th co-ordinate functional (or rather, a coefficient functional) $k_{n}$ is given by
@@
Let $k_{n}$ denote the co-ordinate functional
$f\mapsto\widehat{f}(n),$
defined on $\ell^{p}_{a,b}(\mathtt{A})$ or $c_{0,a,b}(\mathtt{A})$, where $n\in\mathbb{Z}$. Then the following hold.
@@
$(i)$ If $f\in\ell^{p}_{a,b}(\mathtt{A}),$ then $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}(a_{k}+b_{k}z)z^{k}$, and $\|f\|=\left(\sum_{k=-\infty}^{\infty}\lvert\lambda_{k}\rvert^{p}\right)^{1/p}$. Rearrangement into a Laurent series yields $f(z)=\sum_{k=-\infty}^{\infty}(\lambda_{k}a_{k}+\lambda_{k-1}b_{k-1})z^{k},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}.$ It follows, by Hölder inequality, that $|k_{n}(f)|\leq\|f\|_{\ell^{p}_{a,b}(\mathtt{A})}(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}$, and hence $\|k_{n}\|\leq(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}$.
@@
$(iii)$ For $f\in\ell^{p}_{a,b}(\mathtt{A}),$ we have
$(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})(f)$ $=$ $a_{n}f_{n}^{*}(f)+b_{n-1}f_{n-1}^{*}(f)$
$=$ $a_{n}f_{n}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}+b_{n-1}f_{n-1}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}$
@@
Hence, $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\in\mathbb{Z}$. ∎
We recall that the unilateral versions $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ were introduced studied in [15]. These spaces correspond to the normalized bases $(a_{n}+b_{n}z)z^{n}$, where $n\geq 0$. The results in Proposition 2.2 hold for $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ as well.
@@
Let $\mathcal{E}$ be a Banach space of analytic functions on an annulus $\mathtt{A}$ or a disc centered around 00, having evaluation functionals bounded at each $z\in\mathtt{A}$. Then, $k_{n}$ is bounded for all $n\in\mathbb{Z}$. Further, $\text{span}~{}\big{\{}k_{n}:~{}n\in\mathbb{Z}\big{\}}$ is dense in $\mathcal{E}^{*}$ with respect to the weak-topology, and also, in the norm topology of $\mathcal{E}^{*}$ if it is reflexive.
@@
Let $ev_{z}$ denote the evaluation functional acting on $\mathcal{E}$. Since the analyticity of $z\mapsto ev_{z}$ in the strong and norm operator topologies are equivalent, we see that $z\mapsto ev_{z}$ is an $\mathcal{E}^{*}$-valued norm-analytic function. It then admits a norm-convergent Laurent series form $ev_{z}=\sum_{n=-\infty}^{\infty}L_{n}z^{n},\hskip 14.22636ptz\in\mathtt{A}$, for some $L_{n}\in\mathcal{E}^{*}$. For $f\in\mathcal{E}$, we have $f(z)=ev_{z}(f)=\sum_{n=-\infty}^{\infty}L_{n}(f)z^{n}.$ Expanding $f(z)$ in its Laurent series, we must then have $L_{n}=k_{n}$ for all $n$.
To show that $\{k_{n}\}$ is a weak-total set in $\mathcal{E}^{*}$, let $\Lambda\in\mathcal{E}^{**}$ be such that $\Lambda(k_{n})=0$ for all $n$. Hence, there is some $f\in\mathcal{E}$ such that $\lambda(k_{n})=k_{n}(f)=0,~{}\forall n\in\mathbb{Z}$, which shows that all the coefficients of $f(z)$ vanish. Consequently, $f=0$.
We now obtain conditions so that the Laurent monomials $z^{\nu}$ belong to the space $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. The norm estimates, stated below, will be later used in the dynamics of weighted forward shifts.
@@
$\lVert z^{\nu}\rVert\leq M_{1}\left\lvert\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right\rvert~{}~{}~{}\text{and}~{}~{}~{}\lVert z^{-\nu}\rVert\leq M_{2}~{}~{}\left\lvert\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\right\rvert,$ (2.2)
@@
By the basis expansion in $\ell^{p}_{{a},{b}}(\mathtt{A})$ and a rearrangement into Laurent series, we can find some $\{\lambda_{j}\}_{j=-\infty}^{\infty}\in\ell^{p}(\mathbb{Z})$ such that
@@
Then, for a fixed $\nu\geq 2,$
$a_{0}$ $=$ $f_{0}-b_{0}z$
$=$ $f_{0}-\frac{b_{0}}{a_{1}}f_{1}+\frac{b_{0}b_{1}}{a_{1}}z^{2}$
$=$ $f_{0}-\frac{b_{0}}{a_{1}}f_{1}+\frac{b_{0}b_{1}}{a_{1}}f_{2}+\cdots+(-1)^{\nu-1}\frac{b_{0}b_{1}\cdots b_{\nu-2}}{a_{1}a_{2}\cdots a_{\nu-1}}f_{\nu-1}+(-1)^{\nu}\frac{b_{0}b_{1}\cdots b_{\nu-1}}{a_{1}a_{2}\cdots a_{\nu-1}}z^{\nu}.$
It follows that
$z^{\nu}=(-1)^{\nu}\left(\frac{a_{0}a_{1}\cdots a_{\nu-1}}{b_{0}b_{1}\cdots b_{\nu-1}}\right)-(-1)^{\nu}\left(\frac{a_{1}a_{2}\cdots a_{\nu-1}}{b_{0}a_{1}\cdots b_{\nu-1}}f_{0}-\cdots+(-1)^{\nu-1}\frac{1}{b_{\nu-1}}f_{\nu-1}\right).$
Again we can write

### v2

@@ Abstract.
This article aims to initiate a study of the forward weighted shift operator, denoted by $F_{w}$, defined on the Banach spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. Here, $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a.b}(\mathtt{A})$, respectively, are the Banach spaces of analytic functions on a suitable annulus $\mathtt{A}$ in the complex plane, having a normalized Schauder basis of the form $\{(a_{n}+b_{n}z)z^{n},~{}~{}n\in\mathbb{Z}\}$, which is equivalent to the standard basis in $\ell^{p}(\mathbb{Z})$ and $c_{0}(\mathbb{Z})$. We first obtain necessary and sufficient conditions for $F_{w}$ to be bounded, and show that, under certain conditions on $\{a_{n}\}$ and $\{b_{n}\}$, the operator $F_{w}$ is similar to a compact perturbation of a bilateral weighted shift on $\ell^{p}(\mathbb{Z})$. This also allows us to obtain the essential spectrum of $F_{w}$. Further, we study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, supercyclic, and chaotic, and provide a class of chaotic operators that are compact perturbations of weighted shifts on $\ell^{p}(\mathbb{Z})$. Finally, it is proved that the adjoint of a shift on the dual of $\ell^{p}_{a,b}(\mathtt{A})$ can have non-trivial periodic vectors, without being even hypercyclic. Also, the zero-one law of orbital limit points fails for $F_{w}^{*}$.
@@ 1. Introduction
Among various classes of operators that are commonly investigated in operator theory and linear dynamics, the class of weighted shifts has a predominant role. Shift operators are fairly simple to define, yet they yield diverse and appealing results. In operator theory, weighted shifts are well studied in connections with model theory (representation of various operators using shifts), invariant subspaces, cyclic vectors, subnormality, and so on, cf. Halmos [24], and Shields [36]. In the context of linear dynamics it is a usual practice that dynamical properties are first tested for weighted shifts defined on the sequence spaces $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$. Important dynamical notions such as hypercyclicity, chaos, and ergodicity are well distinguishable using weighted shifts. Motivated by these, in this paper we introduce and study weighted shift operators on the spaces $\ell^{p}_{a,b}(\mathbb{D})$ and $\ell^{p}_{a,b}(\mathtt{A})$ having a close resemblance to the weighted $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ spaces, highlighting the similarities and differences in the dynamics. These newly introduced spaces and weighted shifts are of independent interests as well. Our main focus is on the linear dynamics of shifts.
@@
  • topologically transitive if, for two non-empty open sets $U_{1}$ and $U_{2}$ of $X$, there exists a natural number $k$ such that $T^{k}(U_{1})\cap U_{2}\neq\phi$,
@@
Rolewicz [33] showed that $\lambda B$ is hypercyclic on $\ell^{p}(\mathbb{N})$, $1\leq p<\infty$, where $|\lambda|>1$ and $B$ is the unweighted backward shift operator. Also, for particular classes of hypercyclic weighted shifts, see Kitai [25], Gethner and Shapiro [18], and Godefroy and Shapiro [19]. A complete characterization of hypercyclic weighted shifts on $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ was obtained by Salas [34]. In [19] the authors introduced and studied the notion of chaos for operators, and established that several familiar operators including weighted shifts are chaotic. A complete characterization of hypercyclic and chaotic weighted shifts in a general $F$-spaces was obtained by Grosse-Erdmann [22]. We also note that the weighted shifts which are (topologically) mixing were characterized by Costakis and Sambarino [12]. For an account on the theory on hypercyclicity and related topics, we refer to the monographs by F. Bayart and E. Matheron [4], and K.-G. Grosse-Erdmann and A. Peris [23]. Also, see S. Grivaux et al., [21]. In practice, the existence of hypercyclic and chaotic properties of operators is established using well known criteria that are stated below, cf. [4] and [23]. We will make use of these standard criteria to obtain the hypercyclic and chaotic properties of weighted shifts defined on the weighted spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ which are introduced in Section $2$.
@@
(Gethner-Shapiro Criterion [18]) Let $T$ be a bounded operator on a separable Banach space $X$, and let $D$ be a dense subset of $X$. If $\{n_{k}\}\subseteq\mathbb{N}$ is a strictly increasing sequence and $S:D\rightarrow D$ is a map such that, for each $x\in D$, $lim_{k\rightarrow\infty}T^{n_{k}}x=0=\lim_{k\rightarrow\infty}S^{n_{k}}x,$ and $TSx=x$, then $T$ is hypercyclic. In addition, if $n_{k}=k$ for all $k\geq 1$, then $T$ is mixing on $X$.
@@
(Chaoticity Criterion [9]) Let $X$ be a separable Banach space, $D$ be a dense set in $X$, and let $T$ be a bounded operator on $X$. If there exists a map $S:D\rightarrow D$ such that $\sum_{n\geq 0}T^{n}x$ and $\sum_{n\geq 0}S^{n}x$ are unconditionally convergent, and $TSx=x$, for each $x\in D$, then the operator $T$ is chaotic and mixing on $X$.
@@
(Supercyclicity Criterion [31]) For a bounded operator $T$ on a separable Banach space $X$, if there exists a dense subset $D$, a sequence $(n_{k})$ of natural numbers, and a map $S:D\rightarrow D$ such that, for all $x,y\in D$, $\lim_{k\rightarrow\infty}\|T^{n_{k}}x\|\|S^{n_{k}}y\|=0$ and $TSx=x$, then $T$ is supercyclic.
Motivated by the classical bilateral shifts, in this paper we initiate a study of weighted bilateral forward shift operators $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. In Section $2$, we introduce the aforementioned spaces as analytic function spaces and prove that the co-ordinate functionals on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ are continuous. In Section $3,$ we obtain necessary and sufficient conditions for the boundedness of $F_{w}$, and as main results, we study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, mixing, supercyclic, and chaotic. Dynamics of the adjoint of a unilateral forward shift is studied in Section $4$. In Section $5$, it is shown that the bilateral shifts on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ and also the unilateral shifts have distinguishing properties compared to classical weighted shifts . For example, among other results, it is shown that the unweighted unilateral and bilateral forward shifts $F$ such that $F^{*}$ can admit a non-zero periodic vector, without being hypercyclic. It is also shown that the zero-one law of orbital limit points for the adjoint $F^{*}$.
@@
if and only if $\lVert f\rVert_{\ell^{p}_{a,b}(\mathtt{A})}:=\left(\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert^{p}\right)^{\frac{1}{p}}<\infty.$ Note that for $b_{n}=0$ for all $n\in\mathbb{Z},$ we get a space that is equivalent to a weighted $\ell^{p}(\mathbb{Z})$. In a similar way, we define the space $c_{0,a,b}(\mathtt{A})$ which consists of analytic functions, for which $\{(a_{n}+b_{n}z)z^{n}:n\in\mathbb{Z}\}$ forms a normalized Schauder basis, equivalent to the usual basis in $c_{0}(\mathbb{Z})$. Moreover, for $f=\sum_{n}\lambda_{n}f_{n}$ in $c_{0,a,b}(\mathtt{A})$, we have $\lambda_{n}\rightarrow 0$, as $|n|\rightarrow\infty$. Define its norm as
@@
A few comments are in order. The natural domains $\mathtt{A}_{j}$, $j=1,2,3$, will always be assumed to be non-degenerate. Secondly, in the case of the usual weighted spaces, the natural domains of the map $z\mapsto ev_{z}$ are open or closed annuli. However, this is not so for $\ell^{p}_{a,b}(\mathtt{A})$ which is an important property that distinguishes these spaces from the classical weighted $\ell^{p}$ spaces, which will be discussed and utilized in Section $5$. Also, the series $\sum_{n=-\infty}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}$ can still converge at a point $z$ that lies outside $\mathtt{A}_{j}$, and hence the domain of $z\mapsto ev_{z}$ can be extended to “extra points” outside $\mathtt{A_{j}}$, cf. Section $5$. We note that such a phenomenon in the unilateral case when $p=2$, was observed by Adams and McGuire [2] who introduced the space the analytic reproducing kernel Hilbert space $\ell^{2}_{a,b}$ having an orthonormal basis of the form $(a_{n}+b_{n}z)z^{n}$, $n\geq 0$.
Next, we get the continuity of the co-ordinate functionals which are defined as follows. Since each $f(z)$ in $\ell^{p}_{a,b}(\mathtt{A})$ is analytic on the annulus $\mathtt{A}$, we have the Laurent expansion $f(z)=\sum_{n=-\infty}^{\infty}\widehat{f}(n)z^{n}$, $z\in\mathtt{A}$, where $\widehat{f}(n)$ is the $n$-th Laurent coefficient of $f(z)$. The $n$-th co-ordinate functional (or rather, a coefficient functional) $k_{n}$ is given by
@@
Let $k_{n}$ denote the co-ordinate functional $f\mapsto\widehat{f}(n)$, defined on $\ell^{p}_{a,b}(\mathtt{A})$ or $c_{0,a,b}(\mathtt{A})$, where $n\in\mathbb{Z}$. Then the following hold.
@@
$(i)$ If $f\in\ell^{p}_{a,b}(\mathtt{A}),$ then $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}(a_{k}+b_{k}z)z^{k}$, and $\|f\|=\left(\sum_{k=-\infty}^{\infty}\lvert\lambda_{k}\rvert^{p}\right)^{1/p}$. A rearrangement into a Laurent series yields $f(z)=\sum_{k=-\infty}^{\infty}(\lambda_{k}a_{k}+\lambda_{k-1}b_{k-1})z^{k},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}.$ It follows, by Hölder inequality, that
$|k_{n}(f)|\leq\|f\|_{\ell^{p}_{a,b}(\mathtt{A})}(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q},$
and hence $\|k_{n}\|\leq(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}$.
@@
$(iii)$ For $f\in\ell^{p}_{a,b}(\mathtt{A})$ and $n\in\mathbb{Z}$, we have
$(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})(f)$ $=$ $a_{n}f_{n}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}+b_{n-1}f_{n-1}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}$
@@
We recall that the unilateral versions $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ were introduced and studied in [13]. These spaces correspond to the normalized bases $(a_{n}+b_{n}z)z^{n}$, where $n\geq 0$. The results in Proposition 2.2 hold for $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ as well.
@@
Let $\mathcal{E}$ be a Banach space of analytic functions on an annulus $\mathtt{A}$ centered around 00, having evaluation functionals bounded at each $z\in\mathtt{A}$. Then, $k_{n}$ is bounded for all $n\in\mathbb{Z}$. Further,
$\text{span}~{}\big{\{}k_{n}:~{}n\in\mathbb{Z}\big{\}}$
is dense in $\mathcal{E}^{*}$ with respect to the weak-topology, and also, in the norm topology of $\mathcal{E}^{*}$ if it is reflexive.
@@
Let $ev_{z}$ denote the evaluation functional acting on $\mathcal{E}$. Since the analyticity of $z\mapsto ev_{z}$ in the strong and norm operator topologies are equivalent, we see that $z\mapsto ev_{z}$ is an $\mathcal{E}^{*}$-valued norm-analytic function. It then admits a norm-convergent Laurent series: $ev_{z}=\sum_{n=-\infty}^{\infty}L_{n}z^{n},\hskip 14.22636pt(z\in\mathtt{A}),$ for some $L_{n}\in\mathcal{E}^{*}$, $n\in\mathbb{Z}$. On the other hand, if $f\in\mathcal{E}$, we have $f(z)=ev_{z}(f)=\sum_{n=-\infty}^{\infty}L_{n}(f)z^{n}.$ Expanding $f(z)$ in its Laurent series, we must then have $L_{n}=k_{n}$ for all $n$.
To show that $\{k_{n}\}$ is a weak-total set in $\mathcal{E}^{*}$, let $\Lambda\in\mathcal{E}^{**}$ be such that $\Lambda(k_{n})=0$ for all $n$. Hence, there is some $f\in\mathcal{E}$ such that $\Lambda(k_{n})=k_{n}(f)=0,$ $\forall n\in\mathbb{Z}$, which shows that all the coefficients of $f(z)$ vanish. Consequently, $f=0$.
We now obtain conditions so that the Laurent monomials $z^{\nu}$ belong to the space $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. The norm estimates, stated below, will be later used in deducing the dynamical properties of weighted shifts.
@@
$\lVert z^{\nu}\rVert\leq M_{1}\prod_{k=0}^{\nu-1}\left|\frac{a_{k}}{b_{k}}\right|~{}~{}~{}\text{and}~{}~{}~{}\lVert z^{-\nu}\rVert\leq M_{2}~{}~{}\prod_{k=1}^{\nu}\left|\frac{b_{-k}}{a_{-k}}\right|,$ (2.2)
@@
Since $\lim\sup_{n\rightarrow+\infty}\big{|}\frac{b_{n}}{a_{n+1}}\big{|}<1,$ we can find $r<1$ and an integer $N_{1}\geq 1,$ such that $\big{|}\frac{b_{n}}{a_{n+1}}\big{|}<r,$ for all $n\geq N_{1}.$ Also, as $\lim\sup_{n\rightarrow+\infty}\big{|}\frac{a_{-n+1}}{b_{-n}}\big{|}<1$, there exists $R<1$ such that $\left|\frac{a_{-n+1}}{b_{-n}}\right|<R$, for all $n\geq N_{2}$.
The space $\ell^{p}_{a,b}(\mathtt{A})$ contains the constant functions. Indeed, by the basis expansion in $\ell^{p}_{{a},{b}}(\mathtt{A})$ and a rearrangement into Laurent series, we can find some $\{\lambda_{j}\}_{j=-\infty}^{\infty}\in\ell^{p}(\mathbb{Z})$ such that
@@
By the choices of $N_{1}$ and $N_{2}$, we separately obtain the convergence of $\sum_{j>0}|\lambda_{j}|^{p}$ and $\sum_{j<0}|\lambda_{j}|^{p}$, and hence $1\in\ell^{p}_{a,b}(\mathtt{A})$.
For a fixed $\nu\geq 1,$
$a_{0}$ $=$ $f_{0}-b_{0}z=f_{0}-\frac{b_{0}}{a_{1}}f_{1}+\frac{b_{0}b_{1}}{a_{1}}z^{2}$
$=$ $f_{0}+\sum_{j=1}^{\nu-1}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}+(-1)^{\nu}\frac{b_{0}b_{1}\cdots b_{\nu-1}}{a_{1}a_{2}\cdots a_{\nu-1}}z^{\nu}.$
It follows that by using the expansion of the constant function $1$, we obtain
$z^{\nu}=(-1)^{\nu}\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\left(\sum_{j=-\infty}^{\infty}\lambda_{j}f_{j}\right)-(-1)^{\nu}\left(\frac{a_{1}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}f_{0}-\cdots+(-1)^{\nu-1}\frac{1}{b_{\nu-1}}f_{\nu-1}\right).$
Thus, we get
$z^{\nu}$ $=$ $(-1)^{\nu}\frac{1-\lambda_{0}a_{0}}{b_{-1}}~{}\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{\nu}(\lambda_{0}a_{0}-1)\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\left(\frac{1}{a_{0}}f_{0}+\sum_{j=1}^{\nu-1}(-1)^{j}\frac{b_{0}\cdots b_{j-1}}{a_{0}\cdots a_{j}}f_{j}\right)$
$+\frac{\lambda_{0}a_{0}}{a_{\nu}}\left(f_{\nu}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{\nu}b_{\nu+1}\cdots b_{\nu+j-1}}{a_{\nu+1}a_{\nu+2}\cdots a_{\nu+j}}f_{\nu+j}\right),$
which is the expansion of $z^{\nu}$ in terms of the basis $\{f_{j}\}$. Therefore, for $\nu\geq 1$,